# GMODiff: One-Step Gain Map Refinement with Diffusion Priors for HDR Reconstruction

### v1

@@ 1 Introduction
Building on these principles, we propose GMODiff, a gain map-driven one-step diffusion model for multi-exposure HDR reconstruction. Rather than training a LDM to directly reconstruct full HDR content, we adopt a two-stage fine-tuning strategy to adapt a pre-trained multi-step LDM for single-step GM refinement. Specifically, in the first stage, we train a Degradation-aware Regressor Network (DaReg) via a dual-learning strategy to produce an initial GM and extract spatial embeddings that highlight regions where the estimate is unreliable. This joint learning scheme provides the LDM with an informative prior for initialization and directs its attention to ambiguous or error-prone areas, significantly accelerating convergence and improving structural fidelity. In the second stage, we fine-tune the LDM to perform one-step denoising conditioned on these regression-based priors, enabling accurate GM correction and high-fidelity HDR reconstruction at dramatically reduced computational cost. In summary, our main contributions are threefold:
@@ 2.2 Diffusion Models
Diffusion models have emerged as a powerful class of generative frameworks, capable of synthesizing high-fidelity data through iterative denoising of random noise. They have already demonstrated impressive performance across a wide range of low-level vision tasks—such as super-resolution [yue2025arbitrary, chen2025adversarial], denoising [du2025superpc], image enhancement [lan2025efficient], colorization [du2025superpc], and deblurring [xie2025diffusion], thanks to their ability to model fine-grained details and generate visually compelling outputs, often surpassing traditional image restoration methods. Despite their success, the practical deployment of diffusion models remains hindered by their high computational cost, which arises from both complex network architectures [song2020denoising, lu2022dpm]] and the requirement for hundreds of iterative sampling steps during inference. To mitigate this bottleneck, recent efforts have focused on accelerating diffusion by drastically reducing the number of denoising steps. However, aggressive step reduction typically leads to a noticeable degradation in output quality. Recently, img2img [parmar2024one] proposed a general framework that adapts pre-trained single-step diffusion models [xu2024ufogen, wang2023prolificdreamer] to new tasks and domains via adversarial learning. This approach enables efficient inference while leveraging the rich internal knowledge embedded in large-scale pre-trained diffusion models. Nevertheless, directly applying this efficient diffusion paradigm to HDR generation remains challenging. Most pre-trained diffusion models rely on autoencoders trained on 8-bit LDR images, which compress HDR’s high bit-depth into a perceptually encoded, low-dimensional latent space, severely limiting their ability to capture the full dynamic range of the real world.
@@ 3.1 Preliminaries
As shown in Fig. 3 (a), while the variational autoencoder (VAE) of a LDM can encode visually plausible LDR representations that resemble HDR content under standard viewing conditions, it fundamentally fails to preserve true high dynamic range—particularly in highlights and shadows. Moreover, since the denoising network operates in an 8-bit perceptual latent space, a naive solution to enable HDR generation would be to fine-tune both the VAE and the denoising network on HDR data. However, this approach demands large-scale HDR training sets and discards the rich generative priors learned from massive LDR images [wang2025lediff].
@@ 3.2 Degradation-aware Regressor Network
where $\mathbf{c}_{L}$ aggregates encoder features across scales and flattens them into vector representations. The decoder reconstructs the initial gain map using $\mathbf{c}_{L}$, while the segmentation head predicts a binary reliability mask for degraded regions (e.g., due to motion or saturation). This mask endows $\mathbf{c}_{L}$ with awareness of uncertain areas, guiding subsequent refinement.
@@
[…]
@@
where $\mathcal{T}(\cdot)$ denotes the $\mu$-law funtion, $\phi_{i,j}(\cdot)$ is the $j$-th feature map of VGG19 after the $i$-th max-pooling layer, and $\lambda=1\mathrm{e}{-2}$ is a weighting hyperparameter.
Degradation-Aware Embedding. By incorporating a reconstruction objective, DaReg implicitly learns regression-based priors that facilitate dynamic range expansion. To better guide the LDM in refining the initial GM, we explicitly train it to identify a degradation-aware mask, such that $\mathbf{c}_{L}$ implicitly encodes regions where $\hat{\mathbf{G}}$ is unreliable. This guides the refinement process to focus on uncertain regions while avoiding hallucinating implausible details in well-reconstructed areas.
Specifically, given an input LDR sequence, we extract multi-scale latent representations with the DaReg encoder as before. For the degradation-aware mask $M$, we upsample all feature maps to the original resolution, concatenate them, and feed the aggregated features into a $1\times 1$ convolutional layer to produce $\bar{\mathbf{c}}_{L}$. The segmentation head $h(\cdot;\phi)$ processes $\bar{\mathbf{c}}_{L}$ through two convolutional layers with 64 hidden channels followed by a sigmoid activation, outputs a predict mask: $\hat{M}=h(\bar{\mathbf{c}}_{L};\phi),$ where $\hat{M}\in[0,1]$ indicates the predicted degradation at pixel $p$. We obtain the ground-truth mask $M$ using a simple yet effective pixel-consistency criterion:
@@
where $p$ denotes a pixel location, $\|\cdot\|_{1}$ averages the absolute difference across RGB channels to yield a scalar, and $\sigma=4/255$ is a fixed threshold. This formulation is motivated by the observation that static regions yield consistent reconstructions—even with simple image blending—whereas motion or overexposure causes significant pixel-level discrepancies. The loss function adopted is the binary cross-entropy (BCE) loss between $\hat{M}$ and M{M}.
@@ 3.3 Prior-Guided One-Step Diffusion Refinement
Specifically, we obtain multi-scale features $\mathbf{Z}_{L}^{i}$ of the initial GM $\hat{\mathbf{G}}$ via the VAE encoder through forward propagation, while $\mathbf{\hat{Z}}_{H}$ is derived from a one-step diffusion process. Although the initial gain map, reconstructed via a regression approach, endows $\mathbf{Z}_{L}^{i}$ with rich detail fidelity, it may introduce noticeable artifacts. To address this, we propose a Degradation-Aware Refinement Block guided by the embedding $\mathbf{c}_{L}$, which leverages degradation-aware embedings to identify unreliable regions in $\mathbf{Z}_{L}^{i}$. As illustrated in Fig. 2 (c), the Degradation-Aware Refinement Block first concatenates the decoder intermediate features $\mathbf{\hat{Z}}_{H}^{i}$ with $\mathbf{Z}_{L}^{i}$ and applies a $3\times 3$ convolution to obtain refined features $\hat{\mathbf{Z}}^{i}$:
@@
Training Strategy During the training phase, the loss function comprises the GM reconstruction loss and HDR reconstruction loss. Following [xu2024ufogen], we also compute the adversarial loss $\mathcal{L}_{\text{adv}}$ between the reconstructed HDR image and the GT HDR image in the $\mu$-law domain. The total objective for decoder training is:
@@
Additionally, considering that the initial GM $\hat{\mathbf{G}}$ obtained in the first stage may overfit to the training set, which impairs GMODiff’s ability to handle its degradations, we propose a group training strategy for GMODiff to learn robust degradation mitigation. Specifically, we divide the training set into N groups and train $N+1$ DaRegs accordingly, with the extra model trained on the entire training dataset. During GMODiff training, we randomly select one DaReg to provide $\hat{\mathbf{G}}$ and $\mathbf{c}_{L}$. The training objective for GMODiff is ultimately defined as the expected loss over such random selections:
@@ 4 Experiments
Table 1: Quantitative comparison across methods. Higher ($\uparrow$) is better for PU-PSNR,PU-SSIM,MANIQA, MUSIQ, CLIPIQA; lower ($\downarrow$) is better for LPIPS, DISTS, and NIQE. Best and second-best results per metric are highlighted in red and blue, respectively.
Metric DeepHDR [wu2018deep] NHDRR [yan2020deep] AHDRNet [yan2019attention] HDRGAN [niu2021hdr] HDR-Trans [liu2022ghost] SCTNet [tel2023alignment] SAFNet [kong2024safnet] DiffHDR [yan2023toward] AFUNet [li2025afunet] Ours
@@ 4.1 Experimental Settings
Datasets. We collect publicly available real-world datasets commonly used for multi-exposure HDR reconstruction from multiple sources [kalantari2017deep, tel2023alignment, kong2024safnet], resulting in a total of 836 image pairs. Among these, 758 pairs are used for training and the remaining 78 for testing. Each sample consists of three LDR images captured with exposure levels of either {-2, 0, +2} or {-3, 0, +3} stops. Notably, Challenge123 [kong2024safnet] also provides static LDR captures for each scene, enabling the generation of five additional training samples per scene compared to the other datasets.
Implementation Details. We implement our model in PyTorch and leverage the accelerate library for mixed-precision training with FP16. In stage one, we optimize DaReg using the Adam optimizer with a learning rate of 2e-4. The training data is preprocessed by extracting $256\times 256$ patches with a stride of 128, and we use a batch size of 16. DaReg adopts the NAFNet architecture [chen2022simple], with the encoder and decoder comprising [2, 4, 4, 8] and [2, 2, 2, 2] NAFBlocks per stage. The model is trained for 200K iterations on 2 NVIDIA RTX 4090 GPUs. In stage 2, we fine-tune Stable Diffusion 2.1 [rombach2022high] using LoRa with rank 16. DaReg is frozen and serves as a fixed module. We employ the AdamW optimizer with a learning rate of 5e-5, weight decay of 1e-10, and a batch size of 12. This stage is trained for 300K iterations on 4 NVIDIA RTX 4090 GPUs.
Evaluation Metrics.To ensure fair and comprehensive comparison, we employ both perceptual and distortion-based image quality metrics. For full-reference perceptual evaluation, we use LPIPS and DISTS, which assess perceptual similarity via features from pre-trained deep networks. For no-reference assessment, we adopt MANIQA, MUSIQ, and CLIPIQA, all trained to predict human-perceived quality. In particular, as these metrics are not HDR-specific, we computed them in the tone-mapped domain using a standard $\mu$-law operator. For distortion-based metrics, we include PU-PSNR and PU-SSIM, specifically tailored for HDR images, through the official implementation [azimi2021pu21].
@@ 4.2 Comparison with SOTA Methods
Compared Methods. We compare GMODiff against state-of-the-art CNN- and Transformer-based methods specifically designed for multi-exposure HDR image reconstruction, including DeepHDR [wu2018deep], NHDRR [yan2020deep], AHDRNet [yan2019attention], SCTNet [tel2023alignment], and HDR-Trans [liu2022ghost]. In addition, we also evaluate against recent advanced approaches: the GAN-based HDRGAN [niu2021hdr], the diffusion-based DiffHDR [yan2023toward], the optical-flow-based SAFNet [kong2024safnet], and the deep-unfolding-based AFUNet [li2025afunet]. For fair comparison, we retrain all these methods on our collected dataset using their official open-source implementations.
Qualitative Results. We present several challenging cases. As shown in Fig. 4, severe camera motion prevents accurate pixel-wise alignment, leaving only semantic cues available to guide reconstruction. The left case of Fig. 5 involves simultaneous human motion and over-saturation, resulting in significant information loss in the input LDR sequence. In the right example of Fig. 5, although the reference frame contains rich visual content, camera motion introduces misalignment that negatively impacts reconstruction quality. It can be observed that existing methods, which directly learn an end-to-end mapping from LDR to HDR, consistently produce noticeable artifacts or ghosting in the aforementioned challenging regions. Although DiffHDR attempts to tackle the ill-posed nature of HDR reconstruction by leveraging the generative capacity of diffusion models, its limited training data prevents it from synthesizing visually plausible results. In contrast, our proposed GMODiff integrates available input cues with strong image priors from pre-trained LDMs to effectively inpaint artifact-prone regions, while avoiding unrealistic hallucinations, resulting in more accurate and perceptually faithful reconstructions.
@@
Quantitative Results. Quantitative comparisons on publicly available datasets are summarized in Tab. 1. Our method consistently outperforms competing approaches across a diverse set of perceptual metrics. Specifically, it achieves notable gains in both full-reference metrics (e.g., LPIPS and DISTS) and no-reference metrics (e.g., MUSIQ, MANIQA, and CLIPIQA), demonstrating its superior ability to reconstruct visually pleasing and perceptually faithful HDR images. Compared to AFUNet [li2025afunet], the state-of-the-art deep unfold method, our approach attains competitive performance in distortion-based metrics (PU-PSNR and PU-SSIM) while achieving significantly better perceptual quality—highlighting that leveraging regression priors to guide pre-trained latent diffusion models enables an effective trade-off between fidelity and perceptual realism. Moreover, in contrast to diffusion-based alternatives such as DiffHDR [yan2023toward], our method fully exploits the rich image priors embedded in large-scale pre-trained LDMs. DiffHDR, trained only on limited in-domain data, lacks robust generic priors and struggles in complex real-world scenarios.

### v2

# GMODiff: One-Step Gain Map Refinement with Diffusion Priors for Efficient HDR Reconstruction

@@ 1 Introduction
Building on these principles, we propose GMODiff, a gain map-driven one-step diffusion model for multi-exposure HDR reconstruction. To the best of our knowledge, GMODiff is the first LDM-based method designed for multi-exposure HDR reconstruction. Rather than training a LDM to directly reconstruct full HDR content, we adopt a two-stage fine-tuning strategy to adapt a pre-trained multi-step LDM for single-step GM refinement. Specifically, in the first stage, we train a Degradation-aware Regressor Network (DaReg) via a dual-learning strategy to produce an initial GM and extract spatial embeddings that highlight regions where the estimate is unreliable. This learning scheme provides the LDM with an informative prior for initialization and directs its attention to ambiguous or error-prone areas, significantly accelerating convergence and improving structural fidelity. In the second stage, we fine-tune the LDM to perform one-step denoising conditioned on these regression-based priors, enabling accurate GM correction and high-fidelity HDR reconstruction at dramatically reduced computational cost. In summary, our main contributions are threefold:
@@ 2.2 Diffusion Models
Diffusion models have emerged as a powerful class of generative frameworks, capable of synthesizing high-fidelity data through iterative denoising of random noise. They have already demonstrated impressive performance across a wide range of low-level vision tasks-such as super-resolution [yue2025arbitrary, chen2025adversarial], denoising [du2025superpc], image enhancement [lan2025efficient], colorization [du2025superpc], and deblurring [xie2025diffusion], thanks to their ability to model fine-grained details and generate visually compelling outputs, often surpassing traditional image restoration methods. Despite their success, the practical deployment of diffusion models remains hindered by their high computational cost, which arises from both complex network architectures [song2020denoising, lu2022dpm] and the requirement for hundreds of iterative sampling steps during inference. To mitigate this bottleneck, recent efforts have focused on accelerating diffusion by drastically reducing the number of denoising steps. However, aggressive step reduction typically leads to a noticeable degradation in output quality. Recently, img2img [parmar2024one] proposed a general framework that adapts pre-trained single-step diffusion models [xu2024ufogen, wang2023prolificdreamer] to new tasks and domains via adversarial learning. This approach enables efficient inference while leveraging the rich internal knowledge embedded in large-scale pre-trained diffusion models. Nevertheless, directly applying this efficient diffusion paradigm to HDR generation remains challenging. Most pre-trained diffusion models rely on autoencoders trained on 8-bit LDR images, which compress HDR’s high bit-depth into a perceptually encoded, low-dimensional latent space, severely limiting their ability to capture the full dynamic range of the real world.
@@ 3.1 Preliminaries
As shown in Fig. 2 (a), while the variational autoencoder (VAE) of a LDM can encode visually plausible LDR representations that resemble HDR content under standard viewing conditions, it fundamentally fails to preserve true high dynamic range-particularly in highlights and shadows. Moreover, since the denoising network operates in an 8-bit perceptual latent space, a naive solution to enable HDR generation would be to fine-tune both the VAE and the denoising network on HDR data. However, this approach demands large-scale HDR training sets and discards the rich generative priors learned from massive LDR images [wang2025lediff].
@@ 3.2 Degradation-aware Regressor Network
where $\mathbf{c}_{L}$ aggregates encoder features across scales and flattens them into vector representations. The decoder reconstructs the initial gain map using $\mathbf{c}_{L}$, while the segmentation head predicts a binary reliability mask for degraded regions (e.g., due to motion or saturation). This mask will be fused with $\mathbf{c}_{L}$ after the first stage of training, which endows $\mathbf{c}_{L}$ with awareness of uncertain regions and guides the subsequent refinement.
@@
During training, we compute the ground-truth gain map $\mathbf{G}$ using Eq. (1), where $\mathbf{L}$ is the reference frame of the LDR sequence. To enable direct supervision with HDR images and ensure differentiability at zero, we replace the logarithmic compression with the $\mu$-law function. Accordingly, the $\exp_{2}(\cdot)$ in Eq. (1) is substituted by the inverse $\mu$-law mapping: $R(x)=\frac{e^{x\cdot\log(1+\mu)}-1}{\mu}$, with $\mu$=100. We supervise the initial gain map prediction $\hat{\mathbf{G}}$ via an $\mathcal{L}_{1}$ regression loss:
@@
where $\mathcal{T}(\cdot)$ denotes the $\mu$-law function, $\phi_{i,j}(\cdot)$ is the $j$-th feature map of VGG19 after the $i$-th max-pooling layer, and $\lambda=1\mathrm{e}{-2}$ is a weighting hyperparameter.
Degradation-Aware Embedding. By incorporating a reconstruction objective, DaReg implicitly learns regression-based priors that facilitate dynamic range expansion. To better guide the LDM in refining the initial GM, we additionally introduce a segmentation head to predict a degradation-aware mask, which allows the refinement process to focus on uncertain regions and avoids hallucinating implausible details in well-reconstructed areas.
Specifically, given an input LDR sequence, we extract multi-scale latent representations with the DaReg encoder as before. For the degradation-aware mask $M$, we process the encoder features to the same number of channels using $1\times 1$ convolutions, and upsample them to the original resolution for concatenation to produce $\bar{\mathbf{c}}_{L}$. The segmentation head $h(\cdot;\phi)$ processes $\bar{\mathbf{c}}_{L}$ through three convolutional layers with 64 hidden channels, followed by a sigmoid activation, and outputs a predict mask: $\hat{M}=h(\bar{\mathbf{c}}_{L};\phi),$ where $\hat{M}\in[0,1]$ denotes the predicted degradation probability at each pixel $p$. Notably, to prevent the gradient of the segmentation head from affecting the reconstruction accuracy, we perform a detach operation before feature processing. The obtained degradation-aware mask is then compressed through a lightweight CNN network in the next training stage and fused with the original $\mathbf{c}_{L}$ to serve as the conditional guidance for the LDMs.
We obtain the ground-truth mask $M$ using a simple yet effective pixel-consistency criterion:
@@
where $p$ denotes a pixel location, $\|\cdot\|_{1}$ averages the absolute difference across RGB channels to yield a scalar, and $\sigma=4/255$ is a fixed threshold. This formulation is motivated by the observation that static regions yield consistent reconstructions-even with simple image blending-whereas motion or overexposure causes significant pixel-level discrepancies. The loss function adopted is the binary cross-entropy (BCE) loss between $\hat{M}$ and M{M}.
@@ 3.3 Prior-Guided One-Step Diffusion Refinement
Specifically, we obtain multi-scale features $\mathbf{Z}_{L}^{i}$ of the initial GM $\hat{\mathbf{G}}$ via the VAE encoder through forward propagation, while $\mathbf{\hat{Z}}_{H}$ is derived from a one-step diffusion process. Although the initial gain map, reconstructed via a regression approach, endows $\mathbf{Z}_{L}^{i}$ with rich detail fidelity, it may introduce noticeable artifacts. To address this, we propose a Degradation-Aware Refinement Block guided by the embedding $\mathbf{c}_{L}$, which leverages degradation-aware embeddings to identify unreliable regions in $\mathbf{Z}_{L}^{i}$. As illustrated in Fig. 3 (c), the Degradation-Aware Refinement Block first concatenates the decoder intermediate features $\mathbf{\hat{Z}}_{H}^{i}$ with $\mathbf{Z}_{L}^{i}$ and applies a $3\times 3$ convolution to obtain refined features $\hat{\mathbf{Z}}^{i}$:
@@
Training Strategy During the training phase, the loss function comprises the GM reconstruction loss and HDR reconstruction loss. Following [DBLP:journals/corr/abs-2503-06537, guo2025compression, xu2024ufogen], we also compute the adversarial loss $\mathcal{L}_{\text{adv}}$ between the predicted latent gain map $\mathbf{\hat{Z}}_{H}$ and the GT latent gain map $\mathbf{{Z}}_{H}$, where the discriminator consists of a frozen pre-trained denoising UNet encoder and two trainable MLP layers. The total objective for decoder training is:
@@
Additionally, considering that the initial GM $\hat{\mathbf{G}}$ obtained in the first stage may overfit to the training set, which impairs GMODiff’s ability to handle its degradations, we propose a group training strategy for GMODiff to learn robust degradation mitigation. Specifically, we divide the training set into N groups and train $N+1$ DaRegs accordingly, with the extra model trained on the entire training dataset. During GMODiff training, we randomly select one DaReg to provide $\hat{\mathbf{G}}$ and $\mathbf{c}_{L}$. The training objective for GMODiff is ultimately defined as the expected loss over such a random selections:
@@ 4 Experiments
Table 1: Quantitative comparison on the in-distribution test split of our collected dataset. Best and second-best results per metric are highlighted in red and blue, respectively.
Metric DeepHDR [wu2018deep] AHDRNet [yan2019attention] SAFNet [niu2021hdr] HDR-Trans [liu2022ghost] SCTNet [tel2023alignment] AFUNet [li2025afunet] LE-Diff [wang2025lediff] GMDD [guan2025hdr] DiffHDR [yan2023toward] Ours
[…]
Table 2: Quantitative comparisons on the Prabhakar et al.’s [prabhakar2019fast] dataset. Best and second-best results per metric are highlighted in red and blue, respectively.
Metric DeepHDR [wu2018deep] AHDRNet [yan2019attention] SAFNet [niu2021hdr] HDR-Trans [liu2022ghost] SCTNet [tel2023alignment] AFUNet [li2025afunet] LE-Diff [wang2025lediff] GMDD [guan2025hdr] DiffHDR [yan2023toward] Ours
@@ 4.1 Experimental Settings
Datasets. Scenes and motion variations within a single dataset are limited. To better validate the model’s generalization ability under diverse real-world conditions, we collect publicly available real-world datasets [kong2024safnet, tel2023alignment, kalantari2017deep] commonly used for multi-exposure HDR reconstruction, which cover a wide range of illumination and motion conditions. With data augmentation and cropping, we obtain over 400k training pairs in total. For evaluation, we conduct quantitative comparisons on datasets with ground truth, including the test split of our training data (in-distribution) and the Prabhakar [prabhakar2019fast] dataset (out-of-distribution, 562 scenes) for real-world robustness testing. Furthermore, we adopt the datasets of Sen et al. [sen2012robust] and Tursun et al. [Tursun2016data] for qualitative visual assessment only, as they lack ground truth.
Implementation Details. We implement our model in PyTorch and leverage the accelerate library for mixed-precision training with FP16. In stage one, we optimize DaReg using the Adam optimizer with a learning rate of 2e-4. The training data is preprocessed by extracting $256\times 256$ patches, and we use a batch size of 16. DaReg adopts the NAFNet architecture [chen2022simple], with the encoder and decoder comprising [2, 4, 4, 8] and [2, 2, 2, 2] NAFBlocks per stage. The model is trained for 200K iterations on 8 NVIDIA RTX 4090 GPUs. After training, DaReg is frozen, and the segmentation head is independently trained for 5K iterations. In stage 2, we fine-tune Stable Diffusion 2.1 [rombach2022high] using LoRa with rank 16. DaReg and the segmentation head are frozen and serve as fixed modules. We employ the AdamW optimizer with a learning rate of 5e-5, weight decay of 1e-10, and a batch size of 12. This stage is trained for 300K iterations on 4 NVIDIA RTX 4090 GPUs.
Evaluation Metrics. To ensure fair and comprehensive comparison, we employ both perceptual and distortion-based image quality metrics. For full-reference perceptual evaluation, we use DISTS, which assess perceptual similarity via features from pre-trained deep networks. For no-reference assessment, we adopt MANIQA, MUSIQ, and CLIPIQA, all trained to predict human-perceived quality. In particular, as these metrics are not HDR-specific, we computed them in the tone-mapped domain using a standard $\mu$-law operator. For distortion-based metrics, we include PU-PSNR and PU-SSIM, specifically tailored for HDR images, through the official implementation [azimi2021pu21] with a peak luminance of 1000 nits. We compute HDR-VDP2 on linear HDR images with the “rgb-native” color encoding. The data is treated as display-referred, scaled to a peak luminance of 1000 nits, with a viewing angle of $30^{\circ}$ and viewing distance of 0.55 m.
@@ 4.2 Comparison with SOTA Methods
Compared Methods. We compare GMODiff against state-of-the-art CNN- and Transformer-based methods specifically designed for multi-exposure HDR image reconstruction, including DeepHDR [wu2018deep], AHDRNet [yan2019attention], SCTNet [tel2023alignment], and HDR-Trans [liu2022ghost]. In addition, we also evaluate against recent advanced approaches: the diffusion-based DiffHDR [yan2023toward], the optical-flow-based SAFNet [kong2024safnet], and the deep-unfolding-based AFUNet [li2025afunet]. For fair comparison, we retrain all these methods on our collected dataset using their official open-source implementations. We further compare with existing LDM-based HDR reconstruction methods [google2024ultrahdr, guan2025hdr, wang2025lediff], which leverage pre-trained LDMs and are fine-tuned on significantly larger datasets than ours. Notably, since UltraFusion [google2024ultrahdr] is a multi-exposure fusion method that cannot generate HDR images, we only include it for qualitative comparison.
Figure 6: Qualitative comparison of HDR reconstruction results with state-of-the-art LDM-based methods. Our GMODiff outperforms existing LDM-based methods in recovering fine details, preserving physical illumination consistency and suppressing artifacts in overexposed/motion regions.
Qualitative Results. We present several challenging scenarios. As shown in Fig. 4, severe camera motion invalidates accurate pixel-wise alignment, leaving only semantic cues for reconstruction. The left case in Fig. 5 involves both human motion and over-saturation, causing severe information loss in the input LDR sequence. In the right example, despite the informative reference frame, camera motion still induces misalignment and degrades reconstruction quality.
As shown in Fig. 4 and 5, existing end-to-end methods consistently suffer from obvious artifacts or ghosting in these challenging regions. Although DiffHDR exploits the generative ability of diffusion models, it still fails to produce visually plausible results due to limited training data. In contrast, our GMODiff combines input cues with powerful priors from pre-trained LDMs to robustly inpaint artifact-prone regions without unrealistic hallucinations, yielding more accurate and perceptually faithful results. We further compare GMODiff with existing LDM-based methods in Fig.  6. GMDD [guan2025hdr] and LE-Diff [wang2025lediff] exhibit severe detail loss because LDMs operate in a compressed latent space. Specifically, GMDD cannot recover missing content in over-exposed areas, while LE-Diff fails to model physical illumination, leading to large brightness and color deviations from the GT. Ultrafusion [google2024ultrahdr] is trained solely on LDR images and thus cannot generate real HDR outputs; it also relies heavily on optical-flow alignment, producing obvious artifacts when alignment fails. By contrast, guided by the regression module, GMODiff strictly follows physical constraints and avoids latent-compression information loss via the DA Decoder, generating more realistic and visually pleasing results.
Qualitative comparisons with Prabhakar et al. [prabhakar2019fast], Sen et al. [sen2012robust], and Tursun et al. [Tursun2016data] are provided in the supplementary material.
@@
Quantitative Results. Quantitative comparisons on publicly available datasets are summarized in Tab. 1 and Tab. 2. Our method consistently outperforms competing approaches across a diverse set of perceptual metrics. Specifically, it achieves notable gains in both full-reference metrics and no-reference metrics, demonstrating its superior ability to reconstruct visually pleasing and perceptually faithful HDR images. Compared to AFUNet [li2025afunet], the state-of-the-art deep unfold method, our approach attains competitive performance in distortion-based metrics (PU-PSNR and PU-SSIM) while achieving significantly better perceptual quality-highlighting that leveraging regression priors to guide pre-trained latent diffusion models enables an effective trade-off between fidelity and perceptual realism. Moreover, in contrast to diffusion-based alternatives such as DiffHDR [yan2023toward], our method fully exploits the rich image priors embedded in large-scale pre-trained LDMs. DiffHDR, trained only on limited in-domain data, lacks robust generic priors and struggles in complex real-world scenarios. Notably, since LE-Diff [wang2025lediff] recalculates physical brightness during the reconstruction process, its HDR results suffer from severe brightness discrepancies, leading to extremely low full-reference metrics. Therefore, we do not report it in the table. The user study is shown in Figure 9, with details provided in the supplementary material.